OpenStudy (anonymous):

Solve by taking square roots. (If the solution is not a real number, enter NOT REAL in each answer blank.) (y + 5)2 − 50 = 0 y = (smaller value) y = (larger value)

OpenStudy (agreene):

(y + 5)2 − 50 = 0 2y+25-50 = 0 2y-40 = 0 2y=-40 y=20